OpenStudy (anonymous):

Functions: Compute for f(x)=1/x a. f(1/2) b. f(1/10) c. f(0.01) d. f(0.007) Not sure how to do it at all.

OpenStudy (lgbasallote):

f(1/2) just replace x with 1/2

OpenStudy (lgbasallote):

simplify 1/1/2 what do you get?

OpenStudy (anonymous):

Uhh, I'm not sure. Having a really big brain fart. Would it be 1? Or?

OpenStudy (lgbasallote):

\[\huge \frac 1{\frac 12}\] do you know how complex fractions work?

OpenStudy (anonymous):

Sounds Fimilar. But Not 100% Sure.

OpenStudy (lgbasallote):

if you have something like \[\huge \frac 1{\frac 13}\] you just take the reciprocal of 1/3 and multiply it to the numerator \[\huge \implies 1 \times \frac 31\] got it?

OpenStudy (anonymous):

So it would be 2?

OpenStudy (lgbasallote):

yes

OpenStudy (lgbasallote):

f(1/2) = 2

OpenStudy (anonymous):

Okay, How would you do that with decimals?

OpenStudy (lgbasallote):

turn them into fractions

OpenStudy (anonymous):

How would you do that though? A 0.01 would be a 1/100? Or?

OpenStudy (lgbasallote):

yes

OpenStudy (anonymous):

And then a 0.007 would be a 1/700??

OpenStudy (lgbasallote):

nope

OpenStudy (lgbasallote):

7/1000